# pMixFed: Efficient Personalized Federated Learning through Adaptive Layer-Wise Mixup

### v1

@@ 4.1 Mixup
Algorithm 1 pMixFed: Broadcasting global to local model
1:Input: Initial states global model: $GM^{(0)}$, local models: $\{LM^{(0)}_{i}\}_{i=1,\ldots,M}$, Number of communication rounds $T$, number of devices per round $m$, Number of layers in local models $\{L_{i}\}$
@@
4:    Update $\mu$ for each $LM_{i}$, $i=\{1,\ldots,K\}$
5:    Server broadcasts $GM^{(t)}$ to each device in $S(t)$
6:    for each device $m\in S(t)$ in parallel do
7:        $(LM^{(t+1)}_{m},GM^{(t+1)}_{m})=\text{Mixup}[(LM^{\prime(t+1)}_{m},GM^{\prime(t+1)}_{m}),\mu]$
8:        Device sends $GM^{(t+1)}_{m}$ back to server
9:        Update $\mu_{i=1,\ldots,K}$
@@
11:    Server updates $GM^{(t+1)}=\frac{1}{K}\sum_{i\in S(t)}GM^{(t+1)}_{m}$
@@
Algorithm 2 Proposed Mixup for Aggregation
1:Input: Initial states global model: $GM^{(0)}$, Number of communication rounds $T$, Number of local iterations $Itr$, number of devices per round $m$, Mixup degree $\lambda$
@@
3:    Server broadcasts $GM^{(t)}$ to each device in $S(t)$: $LM_{i}^{(t)}$
4:    Update $\lambda_{i=1,\ldots,M}$ for each device
5:    for each device $i\in S(t)$ do
6:        for $epoch=0,1,\ldots,Itr-1$ do
7:           Train local model:
8:                         $LM_{i}^{(t+1)}=GM^{(t)}$
9:                     $GM^{(t+1)}=$
10:$\hskip 65.44133pt\text{Mixup}(LM_{i}^{(t+1)},GM^{(t)})$
11:        end for
12:        Device sends $GM^{(t+1)}$ back to server
13:        Adaptively update $\lambda_{i=1,\ldots,M}$
14:    end for
15:end for
4.2.1 Broadcasting : Global to local model transfer
This stage involves sharing global knowledge with local clients. In the existing PFL methods, the same weight allocation is typically applied to each heterogeneous local model. In our work, we personalize this process by allowing the local model to select the proportion of layers it requires. For instance, for a cold-start user, more information should be extracted from the shared knowledge model, implying that a few layers should be frozen for personalization. Accordingly, A history of the previous global model $G^{t}$ will remain which helps with catastrophic forgetting of the generalized model. Additionally, we introduce a gradual update procedure where the value of $\lambda$ gradually decreases from one (indicating fully shared layers) from the base layer to the end, based on the mixing factor $\mu$. The mix layer is adaptively updated in each communication round for each client individually, according to personalization accuracy. With this adaptive and flexible approach, not only can upcoming streaming unseen data be managed, but also the participation gap (test accuracy of new cold start users) would be improved. The Broadcasting phase is illustrated in 1 which is inspired by [20]. The update rule of local model $L_{k}^{(t)}$ is as follows:
$\small\begin{split}&{L^{\prime}}^{(t)}_{k,i}=(\lambda^{(t)}_{k,i}).G^{(t)}_{i}+(1-\lambda^{(t)}_{k,i}).L^{(t)}_{k,i}\\ &L_{k}^{(t+1)}={L^{\prime}}^{(t)}_{k}-\eta\nabla F_{k}({L^{\prime}}^{(t)}_{k}).\end{split}$ (7)
4.2.2 Aggregation: Local to Global Model Transfer
Existing methods primarily categorize layers into two types: personalized layers and generalized layers. The updated global layers from different clients, $G_{k}^{(t+1)}$, are typically aggregated using Equation 3, which can lead to catastrophic forgetting. Since The base layers of the global model serve as the backbone of shared knowledge [37], This issue arises because, during each local update, the generalized layers undergo substantial modifications [34, 33]. When the global model is updated by simply aggregating local models, valuable information from previously shared knowledge may be lost, leading to forgetting—even if $G^{t}$ performs better than the newly aggregated model. To address this challenge, we propose a new strategy that applies Mixup between the gradients of the previous global model and each client’s local model before aggregation. For each client $i$, the mixup coefficient $\lambda_{i}$ gradually increases from 0 to 1, moving from the head to the base layer, controlled by the mix factor $\mu_{i}$. Additionally, the base layer is adaptively updated based on the communication round and the generalization accuracy, ensuring robust integration of shared and personalized knowledge. It should be noted that parameter $\mu$ is constant in the aggregation stage for all clients as it’s dependent to the average performance of the previous global model.
$\small\begin{split}&{G^{\prime}}^{(t)}_{k,i}=(\lambda^{(t)}_{i,k}).G^{(t)}_{i}+(1-\lambda^{(t)}_{k,i}).L^{(t+1)}_{k,i},\\ &G^{(t+1)}=\sum_{k=1}^{K}\frac{|D_{k}|}{\sum_{k=1}^{K}|D_{k}|}G^{\prime(t+1)}_{k}.\end{split}$ (8)
The high-level block-diagram visualization of the proposed method is shown in Figure 2. It is important to note that the sizes of local models $LM^{(0)}_{i}$ can differ from each other. Consequently, the size of $GM^{(0)}$ should be greater than the maximum size of local models. The parameter $\lambda$ in Equation 4 determines the Mixup degree between the shared model $GM$ and the local models $LM^{(\cdot)}_{i}$, while $\mu$ governs the slope of the change in $\lambda$ across different layers. The degree of Mixup gradually decreases according to the parameter $\mu$ from 00 to $1$. In this scenario, $\lambda=0$ for the first base layer, indicating total sharing, while $\lambda=1$ applies to the final layer, which represents no sharing. The underlying concept is that the base layer contains more general information, whereas the final layers retain client-specific information. The use of the parameter $\mu$, relative to the number of local layers, eliminates the need for a specified cut layer $k$ and allows its application across different model sizes and layers. The parameter $\mu_{i}$ is adaptively updated based on the personalized and global model accuracy for each client. Algorithm 1 shows how Mixup is used as a shared aggregation technique between individual clients and the server. In each training round, only one client is Mixed up with the global model, and $\lambda$ is adaptively learned based on the objective function using online learning. Algorithm 2 shows how Mixup is employed as a shared aggregation technique between the clients and the server. In each training round, only one client is “Mixed up” with the global model and the $\lambda$ parameter is adaptively learned based on the objective function using online learning. For a theoretical convergence analysis of pMixFed, please refer to the Appendix 9.
@@ 5 Experiments
Table 1: Comparison of the State-of-The-Art Methods using Mobile-Net model
@@
Table 2: Comparison of the State-of-The-Art Methods using CNN model
@@ 5.1 Experimental Setup
Datasets: We used three datasets widely used federated learning: MNIST [23], CIFAR-10, and CIFAR-100 [3], We followed the setup in [32] and [33] to simulate heterogeneous, non-IID data distributions across clients for both train and test datasets. The maximum number of classes per user is set to $S=5$ for CIFAR-10 and MNIST, and $S=50$ for CIFAR-100. Experiments were conducted across varying heterogeneous settings, including small-scale ($N=10$) and large-scale ($N=100$) client populations, with different client participation rates $C=[100\%,10\%]$ to measure the effects of stragglers.
@@
Baselines and backbone: We compare two version of our method (i) pMixFed: an adaptive and dynamic mixup-based PFL approach, and (ii) pMixFed-Dynamic: a dynamic-only mixup variant where the parameter $\mu$ is fixed across communication rounds, against several baselines. These baselines include: FedAvg [32], FedAlt [34], FedSim [34], FedBABU [33]. Additionally, we compare against full model personalization methods, Ditto [28], Per-FedAvg [12] , and LG-FedAvg [29]. For all experiments, the number of local training epochs was set to $r=5$, number of communication rounds was fixed at $T=100$, and the batch size was 32. The Adam optimizer has been used while the learning rate, for both global and local updates, was $lr=0.001$ across all communication rounds. The average personalized test accuracy across individual client’s data in the final communication round is reported in Table 2 and 1. Figure 3 presents the training accuracy versus communication rounds for CIFAR10 and CIFAR100 datasets.
@@
Model Architectures: Following the FL literature, we utilized several model architectures. For MNIST, we used a simple CNN consisting of 2 convolutional layers (each with 1 block) and 2 fully connected layers. for CIFAR10 and CIFAR100, We employed a CNN with 4 convolutional layers (1 block each) and 1 fully connected layer for all datasets. Additionally, we used MobileNet, which comprises 14 convolutional layers (2 blocks each) and 1 fully connected layer, for CIFAR-10 and CIFAR-100. For partial model approaches such as FedAlt and FedSim, the split layer is fixed in the middle of the network: for CNNs, layers [1,2] are shared, while for MobileNet, layers [1–7] are considered as shared part. More details about the training process is discussed in Appendix 10.1. Our implementation is available as a supplement for reprodcing the results.
@@
Figure 3: test accuracy accuracy curve along with global communication rounds for pMixFed and PFL baselines experimented on CIFAR10 and CIFAR100 where N=100, C=10%.
@@ 5.2 Comparative Results
We evaluate two versions of our proposed method: pMixFed, which utilizes both adaptive and dynamic layer-wise mixup degree updates, and pMixFed-Dynamic, where the mixup coefficient $\mu_{i}$ is fixed across all training rounds and gradually decreases from $1\rightarrow 0$ from the head to the base layer ($i=M\rightarrow 1$) following a Sigmoid function designed for each client. Our results show that the final accuracy of pMixFed outperforms baseline methods in most cases. As shown in Figure 3, the accuracy curve of pMixFed exhibits smoother and faster convergence, which may suggest the potential for early stopping in FL settings. Previous studies on mixup also suggest that linear interpolation between features provides the most benefit in the early training phase [65]. Moreover, Figure 3 highlights that partial models with a hard split are highly sensitive to hyperparameter selection and different distribution settings, which can lead to training instability. This issue appears to be addressed more effectively in pMixFed, as discussed further in Section 4. According to the results in Table 2 and Table 1, Ditto demonstrates relatively robust performance across different heterogeneity settings. However, its effectiveness diminishes as the model depth increases, such as with MobileNet, and under larger client populations ($N=100$). On the other hand, while FedAlt and FedSim report above-baseline results, they consistently fail during training. Adjusting hyperparameters did not resolve this issue.
@@ 5.4 Ablation Study
Random vs gradual mix factor from $\beta$ distribution: In this paper, we have explored different designs for calculating g mix factor $\mu_{k}$. The value of $\lambda$ in Eq. 4, naturally sampled from a $\beta\neq(\alpha,\alpha)$ distribution [61] which is on the interval [0,1]. We have also experimented the random $\lambda_{i}$ using $\beta$ distribution with different $\alpha$. If $\alpha=1$, the $\beta$ distribution is uniform meaning that the $\lambda$ would be sampled uniformly from [0,1]. Moreover $\alpha>1$, The $\lambda$ would be more in between, creating a more mixed output between $L_{k}$ and $G$. On the contrary, if $\alpha<1$ the mixed model tend to choose just one of the global and local models where $\lambda=1or\lambda=0$. The effects of different $\alpha$ on mixup degree $\lambda$ is discussed in Appendix.
Figure 5: (a) Effect on learning rate on average test accuracy(out-of-sample) gap and on the cold-start users. (b) The comparison between test accuracy on the cold-start-users with different Mix factor functions. (Dynamic-only). In this scenario we used a fixed $Mu$ for all communication rounds. (Sigmoid). The original updating strategy based on a sigmoid function. (Gradual). A simple linear function has been adapted for updating $Mu$. (Random) Mixup degree $\lambda_{i}$ is selected randomly from $\beta$ distribution.
@@
Effect of Mixup Degree as Learning Rate (lr): We observed that the effect of the mixup coefficient is highly influenced by the learning rate and its decay. To empirically demonstrate this relationship, we measured the impact of the learning rate on new participants (cold-start users) as well as on the out-of-sample gap (average test accuracy on unseen data). The results of this comparison are presented in Figure 5 (a). Additional details are provided in the Appendix 10.3.

### v2

@@ 5.1 Mixup
5.2.1 Broadcasting : Global to local model transfer
This stage involves sharing global knowledge with local clients. In the existing PFL methods, the same weight allocation is typically applied to each heterogeneous local model. In our work, we personalize this process by allowing the local model to select the proportion of layers it requires. For instance, for a cold-start user, more information should be extracted from the shared knowledge model, implying that a few layers should be frozen for personalization. Accordingly, A history of the previous global model $G^{t}$ will remain which helps with catastrophic forgetting of the generalized model. Additionally, we introduce a gradual update procedure where the value of $\lambda$ gradually decreases from one (indicating fully shared layers) from the base layer to the end, based on the mixing factor $\mu$. The mix layer is adaptively updated in each communication round for each client individually, according to personalization accuracy. With this adaptive and flexible approach, not only can upcoming streaming unseen data be managed, but also the participation gap (test accuracy of new cold start users) would be improved. The Broadcasting phase is illustrated in Figure 1 which is inspired by kairouz2021advances. The update rule of local model $L_{k}^{(t)}$ is as follows:
[…]
5.2.2 Aggregation: Local to Global Model Transfer
Existing methods primarily categorize layers into two types: personalized layers and generalized layers. The updated global layers from different clients, $G_{k}^{(t+1)}$, are typically aggregated using Equation 3, which can lead to catastrophic forgetting. Since The base layers of the global model serve as the backbone of shared knowledge raghu2019rapid, This issue arises because, during each local update, the generalized layers undergo substantial modifications pillutla2022federated; oh2021fedbabu. When the global model is updated by simply aggregating local models, valuable information from previously shared knowledge may be lost, leading to forgetting—even if $G^{t}$ performs better than the newly aggregated model. To address this challenge, we propose a new strategy that applies Mixup between the gradients of the previous global model and each client’s local model before aggregation. For each client $i$, the mixup coefficient $\lambda_{i}$ gradually increases from 0 to 1, moving from the head to the base layer, controlled by the mix factor $\mu_{i}$. Additionally, the base layer is adaptively updated based on the communication round and the generalization accuracy, ensuring robust integration of shared and personalized knowledge. It should be noted that parameter $\mu$ is constant in the aggregation stage for all clients as it’s dependent to the average performance of the previous global model.
[…]
The high-level block-diagram visualization of the proposed method is shown in Figure 2. It is important to note that the sizes of local models $LM^{(0)}_{i}$ can differ from each other. Consequently, the size of $GM^{(0)}$ should be greater than the maximum size of local models. The parameter $\lambda$ in Equation 9 determines the Mixup degree between the shared model $GM$ and the local models $LM^{(\cdot)}_{i}$, while $\mu$ governs the slope of the change in $\lambda$ across different layers. The degree of Mixup gradually decreases according to the parameter $\mu$ from 0 to $1$. In this scenario, $\lambda=0$ for the first base layer, indicating total sharing, while $\lambda=1$ applies to the final layer, which represents no sharing. The underlying concept is that the base layer contains more general information, whereas the final layers retain client-specific information. The use of the parameter $\mu$, relative to the number of local layers, eliminates the need for a specified cut layer $k$ and allows its application across different model sizes and layers. The parameter $\mu_{i}$ is adaptively updated based on the personalized and global model accuracy for each client. Algorithm 1 shows how Mixup is used as a shared aggregation technique between individual clients and the server. In each training round, only one client is Mixed up with the global model, and $\lambda$ is adaptively learned based on the objective function using online learning. Algorithm 2 shows how Mixup is employed as a shared aggregation technique between the clients and the server. In each training round, only one client is “Mixed up” with the global model and the $\lambda$ parameter is adaptively learned based on the objective function using online learning.
Algorithm 1 Broadcasting: global to local model transfer
1:Input: Initial states global model: $G^{(0)}$, local models: $\{L^{(0)}_{i}\}_{i=1,\ldots,N}$, Number of communication rounds $T$, Number of communication rounds $T$, Number of local training rounds ItrItr, Number of layers in local models $\{L_{i}\}$
@@ 5.2.2 Aggregation: Local to Global Model Transfer
4:  Update $\mu$ for each $L_{i}$, $i=\{1,\ldots,K\}$
5:  Server broadcasts $G^{(t)}$ to each device in $S(t)$
6:  for each device $k\in S(t)$ do
7:   $L^{\prime(t)}_{k}=\text{Mix}[(L^{(t)}_{k},G^{(t)}_{k}),\mu_{Broad}]$
8:   for $l=0,1,\ldots,Itr-1$ do
9:     $L^{\prime(t+1)}_{k}\xleftarrow{Localtrain}(L^{\prime(t)}_{k},\eta)$
@@
11:   Each device saves the updated model $L^{(t+1)}_{k}=L^{\prime(t+1)}_{k}$
@@
13:end for
Algorithm 2 Aggregation:local to global model transfer
1:Input: Initial states global model: $G^{(0)}$, Number of communication rounds $T$, number of devices per round $s(t)$, Mix factor $\mu$
@@
3:  for each device $k\in S(t)$ do
4:   Update $\mu^{t}_{k}$ for $k$ in round $t$
5:   $G^{\prime(t+1)}_{k}=\text{Mix}[(L^{(t+1)}_{k},G^{(t)}_{k}),\mu_{Agg}]$
6:  end for
7:  Each Device sends $G^{\prime(t+1)}_{k}$ back to server
8:  $G^{(t+1)}\xleftarrow{Agg}{G^{\prime(t+1)}_{k}}_{k\in S(t)}$
9:end for
@@ 6 Experiments
Table 2: Average Test Accuracy of different methods using Mobile-Net model. See more details in section 6.2
@@
Table 3: Average Test Accuracy of different methods using CNN model. See more details in section 6.2
@@ 6.1 Experimental Setup
6.1.1 Datasets:
We used three datasets widely used federated learning: MNIST lecun1998gradient, CIFAR-10, CIFAR-100 alex2009learning, and Caltech-101. CIFAR-10 consists of 50,000 images of size $32\times 32$ for training and 10,000 images for testing. CIFAR-100, on the other hand, consists of 100 classes, with 500 $32\times 32$ images per class for training and 100 images per class for testing. MNIST contains 10 labels and includes 60,000 samples of $28\times 28$ grayscale images for training and 10,000 for testing. Caltech-101 is a canonical benchmark for object recognition, introduced by Li and colleagues in 2003. The dataset comprises 9,146 images distributed across 101 object categories—such as helicopter, elephant, and chair—plus a background/clutter class. Category sizes range from roughly 40 to 800 images (with most near 50), and images are typically around 300×200 pixels.
@@ 6.1.1 Datasets:
We followed the setup in mcmahan2017communication and oh2021fedbabu to simulate heterogeneous, non-IID data distributions across clients for both train and test datasets. The maximum number of classes per user is set to $S=5$ for CIFAR-10 and MNIST, and $S=50$ for CIFAR-100. Experiments were conducted across varying heterogeneous settings, including small-scale ($N=10$) and large-scale ($N=100$) client populations, with different client participation rates $C=[100\%,10\%]$ to measure the effects of stragglers. For Caltech-101 experiment setup, please refer to the appendix 10.3.
6.1.2 Training Details:
For evaluation, we have reported the average test accuracy of the global model yuan2021we for different approaches. The final global model at the last communication round is saved and used during the evaluation. The global model is then personalized according to each baseline’s personalization or fine-tuning algorithm for $r=4$ local epochs and $T=50$. For FedAlt, the local model is reconstructed from the global model and fine-tuned on the test data. For FedSim, both the global and local models are fine-tuned partially but simultaneously. In the case of FedBABU, the head (fully connected layers) remains frozen during local training, while the body is updated. Since we could not directly apply pFedHN in our platform setting, we adapted their method using the same hyper parameters discussed above and employed hidden layers with 100 units for the hypernetwork and 16 kernels. The local update process for LG-FedAvg, FedAvg, and Per-FedAvg simply involves updating all layers jointly during the fine-tuning process. The global learning rate for the methods that need sgd update in the global server e.g., FedAvg, has been set from $lr_{global}=[1e-3,1e-4,and1e-5]$.It should be noted that due to the performance drop for some methods (FedAlt , FedSim) in round 10 or 40 in some settings, we’ve reported the highest accuracy achieved. Also this is the reason the accuracy curves are illustrated for 39 rounds instead of 50.  [^4] Change of hyper-parameters such as lr, batch size, momentum and even changing the optimizer to adam didn’t help with the performance drop in most cases.
6.1.3 Baselines and backbone:
@@
We compare our method pMixFed: an adaptive and dynamic mixup-based PFL approach, against several baselines. These baselines include: FedAvg mcmahan2017communication, FedAlt pillutla2022federated, FedSim pillutla2022federated, FedBABU oh2021fedbabu. Additionally, we compare against full model personalization methods, Ditto li2021ditto, FedRep yang2023fedrep Per-FedAvg fallah2020personalized , and LG-FedAvg liang2020think. For all experiments, the number of local training epochs was set to $r=4$, number of communication rounds was fixed at $T=50$, and the batch size was 32. The Adam optimizer has been used while the learning rate, for both global and local updates, was $lr=0.001$ across all communication rounds. The average personalized test accuracy across individual client’s data in the final communication round is reported in Table 3 and 2. Figure 3 presents the training accuracy versus communication rounds for CIFAR10 and CIFAR100 datasets.
6.1.4 Model Architectures:
@@
Following the FL literature, we utilized several model architectures. For MNIST, we used a simple CNN consisting of 2 convolutional layers (each with 1 block) and 2 fully connected layers. for CIFAR10 and CIFAR100, we employed a CNN with 4 convolutional layers (1 block each) and 1 fully connected layer for all datasets. Additionally, we used MobileNet, which comprises 14 convolutional layers (2 blocks each) and 1 fully connected layer, for CIFAR-10 and CIFAR-100. For partial model approaches such as FedAlt and FedSim, the split layer is fixed in the middle of the network: for CNNs, layers [1,2] are shared, while for MobileNet, layers [1–7] are considered as shared part. More details about the training process are discussed in Appendix Sec. 4.1. Experimental Setup. Our implementation is also available as a supplement for reproducing the results.
@@
Figure 3: Average Test accuracy in different global communication rounds for pMixFed and other PFL baselines experimented on CIFAR10 and CIFAR100 where N=100, C=10%. More details are discussed in Section 6.3
@@ 6.2 Comparative Results
We evaluated our proposed method: pMixFed, which utilizes both adaptive and dynamic layer-wise mixup degree updates, and pMixFed-Dynamic, where the mixup coefficient $\mu_{i}$ is fixed across all training rounds and gradually decreases from $1\rightarrow 0$ from the head to the base layer ($i=M\rightarrow 1$) following a Sigmoid function designed for each client. Our results show that the average test accuracy of pMixFed outperforms baseline methods in most cases. As shown in Figure 3, the accuracy curve of pMixFed exhibits smoother and faster convergence, which may suggest the potential for early stopping in FL settings. Previous studies on mixup also suggest that linear interpolation between features provides the most benefit in the early training phase zou2023benefits. Moreover, Figure 3 highlights that partial models with a hard split are highly sensitive to hyperparameter selection and different distribution settings, which can lead to training instability. This issue appears to be addressed more effectively in pMixFed, as discussed further in Section 4. According to the results in Table 3 and Table 2, Ditto demonstrates relatively robust performance across different heterogeneity settings. However, its effectiveness diminishes as the model depth increases, such as with MobileNet, and under larger client populations ($N=100$). On the other hand, while FedAlt and FedSim report above-baseline results, they consistently fail during training.  [^5] Adjusting hyperparameters did not resolve this issue.
Caltech-101. To further evaluate robustness under high inter-class variability and limited per-class samples, we conduct experiments on Caltech-101, a canonical benchmark for heterogeneous federated learning. This dataset poses unique challenges due to its long-tail distribution (most categories have $\sim$50 images) and high intra-class variability. We follow the same federated setup as in CIFAR/MNIST experiments, splitting the data among $N=\{10,100\}$ clients with participation rates $C=\{10\%,100\%\}$. As in prior work, each client is restricted to at most $S=30$ classes. We compare pMixFed against two recent state-of-the-art personalization baselines: FedRep yang2023fedrep, which learns shared representations with personalized heads, and Per-FedAvg fallah2020personalized, which meta-learns a global initialization for fast personalization. We also include FedAlt pillutla2022federated as a representative partial-sharing baseline.
Table 4: Caltech-101 top-1 accuracy (%) with a CNN backbone across four federated settings. Values are the average across evaluation angles A–D; full A–D breakdowns are reported in Appendix 10.3.
Method $N{=}100,\penalty 10000\ C{=}100\%$ $N{=}10,\penalty 10000\ C{=}100\%$ $N{=}100,\penalty 10000\ C{=}10\%$ $N{=}10,\penalty 10000\ C{=}10\%$
pMixFed (Ours) 79.3 82.4 76.1 79.0
FedRep 77.0 80.1 74.2 76.9
Per-FedAvg 76.5 81.0 73.5 76.2
FedAlt 73.1 75.5 69.5 72.2
Across all settings, pMixFed consistently achieves the highest accuracy (Table 4). In particular, under the most heterogeneous regime (N=100N{=}100, $C{=}10\%$), pMixFed improves by $\sim$2% over FedRep and $\sim$3% over Per-FedAvg, while FedAlt lags by 6–7%. Per-FedAvg performs competitively in the low-client regime ($N=10$), but its advantage diminishes as heterogeneity increases. These results validate that pMixFed not only excels on CIFAR and MNIST benchmarks, but also generalizes robustly to harder, imbalanced datasets like Caltech-101.
@@ 6.4 Ablation Study
Random vs gradual mix factor from $\beta$ distribution: In this paper, we have explored different designs for calculating g mix factor $\mu_{k}$. The value of $\lambda$ in Eq. 9, naturally sampled from a $\beta\neq(\alpha,\alpha)$ distribution zhang2017mixup which is on the interval [0,1]. We have also experimented the random $\lambda_{i}$ using $\beta$ distribution with different $\alpha$. If $\alpha=1$, the $\beta$ distribution is uniform meaning that the $\lambda$ would be sampled uniformly from [0,1]. Moreover $\alpha>1$, The $\lambda$ would be more in between, creating a more mixed output between $L_{k}$ and $G$. On the contrary, if $\alpha<1$ the mixed model tend to choose just one of the global and local models where $\lambda=1or\lambda=0$. The effects of different $\alpha$ on mixup degree $\lambda$ is discussed in Appendix sec. 4.3
Figure 5: (a) Effect on learning rate on average test accuracy(out-of-sample) gap and on the cold-start users. (b) The comparison between test accuracy on the cold-start-users with different Mix factor functions. (Dynamic-only). In this scenario we used a fixed MuMu for all communication rounds. (Sigmoid). The original updating strategy based on a sigmoid function. (Gradual). A simple linear function has been adapted for updating MuMu. (Random) Mixup degree $\lambda_{i}$ is selected randomly from $\beta$ distribution.
@@
Effect of Mixup Degree as Learning Rate (lr): We observed that the effect of the mixup coefficient is highly influenced by the learning rate and its decay. To empirically demonstrate this relationship, we measured the impact of the learning rate on new participants (cold-start users) as well as on the out-of-sample gap (average test accuracy on unseen data). The results of this comparison are presented in Figure 5 (a). Additional details are provided in the Appendix sec. 3.4.1. Analytical analysis of the Effect of learning rate and mixup degree.
6.5 Communication and Computational Cost
Although pMixFed introduces additional computation compared to partial personalization baselines (e.g., FedAlt, FedSim), we analyze its efficiency in terms of communication and runtime. Figure 6(a) shows the number of layers frozen at each communication round due to a zero mixup degree, which naturally reduces the amount of information exchanged. Figure 6(b) reports the average communication overhead across 100 clients for CNN and MobileNet on CIFAR-100.
Figure 6: (a) Number of frozen layers per communication round due to zero mixup degree. (b) Average communication overhead per round across 100 clients for CNN and MobileNet on CIFAR-100.
Compared to full-model FL (FedAvg), pMixFed does not increase the parameter size transmitted per layer. While its per-round communication cost is higher than partial methods (FedAlt, FedSim), two key properties compensate for this overhead:
Faster convergence and early stopping. pMixFed reaches target accuracy in significantly fewer communication rounds, reducing the overall cost. In large-scale and highly heterogeneous settings (e.g., MobileNet on CIFAR-100), FedAlt and FedSim often diverge or require far more rounds to achieve comparable accuracy.
Adaptive layer freezing. The mixup degree $\lambda_{i}^{(t)}$, updated dynamically via Eq. (5)–(6), drives progressive freezing of layers. As training advances, an increasing number of layers are excluded from both broadcasting and aggregation, further reducing transferred parameters in later rounds. We also evaluated computational efficiency by measuring GPU usage and runtime per round (Table 5). In the most demanding case (CIFAR-100, MobileNet, $C=100\%$), pMixFed is only slightly slower (1–7s per round) and requires marginally more GPU memory (1–22MB). Given its faster convergence (on average, $\sim$10% fewer rounds) and improved stability, the effective overhead is substantially mitigated.
Table 5: GPU usage and computational time per global round using MobileNet on CIFAR-100 ($C=100\%$).
Method $N=100$ $N=10$
Time (s) Memory (MB) Time (s) Memory (MB)
FedAvg 94.16 5362.44 95.31 735.25
FedAlt 94.75 5372.77 92.86 746.85
FedSim 95.72 5378.19 91.43 746.85
pMixFed 100.03 5376.90 98.48 762.99
In summary, while pMixFed incurs a minor per-round overhead, its faster convergence and adaptive freezing mechanism reduce the overall communication and computation cost, making it more scalable and robust in heterogeneous FL settings.